# Phase diagram and universal scaling regimes of two-dimensional exciton–polariton Bose–Einstein condensates

### v1

@@ IV.7 Distribution of condensate phase and density fluctuations
where $\Delta t\in\left[35{\rm ps},350{\rm ps}\right]$ as shown on Fig. 3, is displayed on Fig. 9. The $\text{sign}(\lambda)$ in (18) stems for the negative value obtained for $\lambda$ with our parameters, which flips the distribution into its mirror image compared with the theoretical predictions obtained for $\lambda>0$ [22, 24]. The distribution of phase fluctuations exhibits clear non-Gaussian tails. As $\Delta t/t_{0}\sim 0.1$ in our simulations, we expect it to be described by the Pearson distribution. The latter is represented with a dashed blue line together with our numerical datas on Fig. 9, showing a perfect agreement with the expected KPZ distribution.
@@
Figure 10: Histograms of the density $n$ normalized by its mean-field value $n_{0}$ corresponding to the EW, KPZ, and vortex points of Fig. 2. Each histogram contains approximately $5\times 10^{5}$ datas. The two formers are fitted by a Gumbel distribution of parameter $m=10$, mean $\langle n\rangle/n_{0}$, and variance $\text{Var}\left[n/n_{0}\right]$, shown by the dashed blue line. The latter is fitted by an exponential distribution of parameter $p=3.5$.
First we observe that the density distribution at the EW and at the KPZ point perfectly match, which is not too surprising from Figs. 2 and 4. However, unexpectedly, these distributions exhibit clear deviations from a Gaussian distribution and have finite skewness and kurtosis. A possible origin could be a subleading coupling between density and phase, but a definite conclusion requires further studies which are beyond the scope of this work. We find that they are well fitted by a Gumbel distribution of parameter $m=10$. At the vortex point, the density distribution markedly differs and is rather fitted by an exponential distribution. This indicates again that the formation of defects occurs randomly, following a Poisson process. Similar exponential distributions have been reported in the nucleation time distribution of spiral defects, in the context of the 2d CGLE [100], and the discrepancies on Fig. 10 between the histogram and the exponential distribution at large $n/n_{0}$ can be explained by finite size effects [70].
@@ V Conclusion
In this paper, through extensive numerical simulations of a generalized Gross-Pitaevskii equation, we have explored the universal properties of a 2d exciton-polariton condensate under incoherent pumping. Controlling the effective nonlinearity of the phase dynamics through the interaction strength between reservoir excitons and polaritons, we have identified three different regimes: a first one for small nonlinearity where the dynamics of the density and of the phase are decoupled, and where the phase exhibits EW universal properties, a second one at intermediate nonlinearity where the density and phase are still decoupled but where the phase roughens with properties belonging to the KPZ universality class, a third one for large values of the nonlinear parameter, where density and phase fluctuations are coupled, since spatial vortices induce both phase singularities and deep minima in the condensate density at the positions of the vortex cores.
@@
We have characterized the properties of these three regimes by studying the condensate first-order correlation function, the momentum distribution, and the distribution of the phase and density fluctuations. In the weakly nonlinear limit, we report evidence of an algebraic decay of the first-order correlation function indicating the emergence of the EW universality class, from which we establish, upon increasing the interaction strength a slow crossover towards stretched exponential decay with exponents of the 2d KPZ class, and eventually a very rapid decay leading to saturation in the vortex phase. This observed behaviors in the EW and KPZ regimes are then confirmed by a careful study of finite size effects in the condensate decoherence time. The different phases are also distinguished by their momentum distribution, where notably the EW and KPZ regimes exhibit different scaling at small momenta, in contrast with 1d condensates.
@@
Finally, we have focused on the statistical properties of phase fluctuations to provide a further validation of the three identified regimes. We have found that the distribution of phase fluctuations is Gaussian in the EW weakly nonlinear regime, it displays non-Gaussian tails at stronger non-linearity in the KPZ regime, in agreement with those predicted by numerical simulations of interfaces in the 2d KPZ class, whereas the distribution is exponential in the vortex regime. Unexpectedly, we observe that the probability distribution of the density fluctuations is not Gaussian in the EW and KPZ regimes, whose explanation deserves further studies.
@@
Our analysis highlights that earlier observations from experiments [84] and numerical simulations [34] correspond to the realization of the EW regime and provides a microscopic prediction for the observed power-law decays and the ratio of the corresponding exponents.
@@
Our work provides a unified picture of the universal properties of 2d polariton condensates in various parameter regimes. In outlook, from the theoretical point of view, it would be of great interest to provide a quantitative prediction for the transition to the vortex regime, whose qualitative picture has been developed in Refs. [30, 31], as well as a deeper understanding of oscillations observed in density correlations. A theoretical explanation for the observed non-Gaussian statistics of the density fluctuations in the EW and KPZ regimes is also missing. Overall, our findings pave the way for further experimental explorations of the universal properties of driven-dissipative quantum systems.
@@ Acknowledgements.
We warmly thank Ivan Amelio, Marzena Szymańska, Harvey Weinberger and Vincent Rossetto for stimulating discussions. FH acknowledges support from the Laboratoire d’excellence LANEF in Grenoble (ANR-10-LABX-51-01), AM acknowledges funding from the Quantum-SOPHA ANR Project ANR-21-CE47-0009. LC acknowledges support from Institut Universitaire de France (IUF). DPD, QF, SR, and JB acknowledge support from the Paris Ile-de-France Region in the framework of DIM QuanTIP, the French RENATECH network, the H2020- FETFLAG project PhoQus (820392), the European Research Council via projects StG ARQADIA (949730) and AdG ANAPOLIS (101054448)
@@ Appendix A Adiabatic Bogoliubov theory
In the adiabatic approximation of the reservoir $\partial_{t}n_{R}\approx 0$, Eq. (20) reduces to $n_{R}=P/(\gamma_{R}+R|\psi|^{2})$. The condensate wavefunction is expanded in fluctuations $\delta\psi$ around its stationary mean field density $n_{0}=\gamma_{R}(p-1)/R$, which reads $\psi(\bm{r},t)=e^{-i\Omega_{0}t}\left[\sqrt{n_{0}}+\delta\psi(\bm{r},t)\right]$. Injecting it in Eq. (19) leads at zeroth order to the mean field exciton reservoir density and condensate blueshift, respectively $n_{R0}=\gamma_{0}/R$ and $\Omega_{0}=gn_{0}+2g_{R}n_{R0}$. To first order in $\delta\psi$, Eq. (19) reduces to the linear dynamics
@@
where we introduced $\tilde{\xi}=\xi e^{-i\Omega_{0}t}$ and the Bogoliubov matrix
@@
$\left(\partial_{t}-\mathcal{F}^{-1}\left[\mathcal{L}_{rot}\right]\right)\begin{pmatrix}\delta n/2n_{0}\\ \delta\theta\end{pmatrix}=\sqrt{\dfrac{\sigma}{n_{0}}}\begin{pmatrix}{\rm Im}\left(\tilde{\xi}\right)\\ -{\rm Re}\left(\tilde{\xi}\right)\end{pmatrix}\,,$ (25)
@@
$\omega_{\pm}=i\Gamma_{\hat{\bm{k}}}\pm\sqrt{E_{\hat{\bm{k}}}^{2}-g_{i}^{2}n_{0}^{2}}\,,$ (27)
@@ A.1 Phase-phase correlator
$+\dfrac{\langle\delta n_{\bm{k},\omega}\delta n_{-\bm{k},-\omega}\rangle}{4n_{0}^{2}}.$
The large scale properties of the condensate can be understood taking advantage of the splitting between these two branches at the crossover momentum $k_{c}=\sqrt{2n_{0}|m|}\left[\sqrt{g_{e}^{2}+g_{i}^{2}}-|g_{e}|\right]^{1/2}$, separating the infrared collective modes $k\ll k_{c}$ from the ultraviolet free particles $k\gg k_{c}$. Focusing on the former limit and choosing $\epsilon_{\hat{\bm{k}}}=k^{2}/2m$ and $\gamma_{\ell}(\hat{\bm{k}})=\gamma_{0}+\gamma_{2}k^{2}$, we find for $k\approx 0$ that $\omega_{+}\approx 2ig_{i}n_{0}$ and $\omega_{-}\approx i\nu k^{2}$, where $\nu$ is the effective phase viscosity (4) introduced in the main text.
@@
×$\times\dfrac{k^{2}}{\left(\omega^{2}+4g_{i}^{2}n_{0}^{2}\right)\left(\omega^{2}+\nu^{2}k^{4}\right)}$
[…]
@@
where $\mathcal{C}_{0}=\frac{D}{\nu}(4\nu)^{2\beta}$ and $y_{0}=2\sqrt{\nu}$ are non-universal constants, with $D$ and $\nu$ the effective parameters defined in Eqs. (4) and (6). This function is characterized by the universal dynamical exponent $z=2$, the growth exponent $\beta=\frac{2-d}{4}$, and the universal scaling function $F^{(d)}(y)=\int\frac{d\bm{k}}{(2\pi)^{d}}\frac{1-e^{i\bm{k}\cdot\bm{y}-\bm{k}^{2}/4}}{\bm{k}^{2}}$ [101, 44].
Let us analyze the finite-size effects on $C_{\rm EW}(r=0,t)$. For a system of linear size $L$, from the diffusive nature of EW modes, one expects a change in its behavior at a time $t^{*}\sim L^{2}/\nu$. Following the arguments given in Ref. [82], we can split the dimensionless momentum integral of $F^{(d)}(y)$ into two: one for $k^{2}<4$ (where $e^{-k^{2}/4}$ is expanded to second order), the other for $k^{2}>4$ (where $e^{-k^{2}/4}\approx 0$). Reintroducing the dimensions, we can then take the short or long time limits of each of them, from which we obtain:
$\bullet$ in dimension d = 1 : $C_{\rm EW}(0,t\ll t^{*})\approx\dfrac{2D}{\pi}\sqrt{t/\nu}$ consistent with the scale invariant form (29). In the limit $t\gg t^{*}$, $C_{\rm EW}(0,t)\approx\dfrac{4D}{\mathcal{L}}t$ where $\mathcal{L}$ is a length scale maintaining homogeneity. A natural choice for this length scale, stemming from the lowest momenta of the finite-size system, is $\mathcal{L}=\uptau^{(1)}L$, with $\uptau^{(1)}$ a dimensionless constant [86]. This constant contains the dependence of the condensate linewidth on the microscopic parameters that goes beyond the Bogoliubov approximation.
$\bullet$ in dimension d = 2: Similarly we obtain for $t\ll t^{*}$ that $C_{\rm EW}(0,t)\approx\dfrac{D}{2\pi\nu}\left[1+\ln(\nu t\Lambda^{2})\right]$, i.e. a logarithmic time dependence [64, 65, 102, 44, 55] with $\Lambda$ a momentum cutoff. When $t\gg t^{*}$, $C_{\rm EW}(0,t)\approx\dfrac{D}{\uptau^{(2)}L^{2}}t$ with the dimensionless constant $\uptau^{(2)}$.
@@
From this finite-size study, we expect $g^{(1)}(t\gg t^{*})\sim e^{-t/\uptau_{c}(L)}$ with the coherence time $\uptau_{c}(L)\propto L^{z-2\beta z}$, generalizing the result of [103, 86] to 2d systems. This Schawlow-Townes-like limit [87], following EW or KPZ features, means that the growing phase interface looses the memory of its initial condition at long times and saturates to a Brownian motion [88].
@@ A.2 Momentum distribution
$n_{\hat{\bm{k}}}=2\pi n_{0}\delta(\bm{k})+\dfrac{\sigma}{\Gamma_{\hat{\bm{k}}}}\left[1+4n_{0}^{2}\dfrac{g_{e}^{2}+g_{i}^{2}}{E_{\hat{\bm{k}}}^{2}+\Gamma_{\hat{\bm{k}}}^{2}-g_{i}^{2}n_{0}^{2}}\right].$ (32)
[…]
At equal time, we obtain $\tilde{C}_{\rm EW}(k,0)\sim k^{-2\beta z-d}$. From this, we deduce that in the EW regime $n_{\hat{\bm{k}}}\propto k^{-2}$. Since this result was obtained using the linear theory, we can expect the non-linearity of the phase dynamics to affect this regime and rather lead to $n_{\hat{\bm{k}}}\underset{k\to 0^{+}}{\sim}\tilde{C}_{\rm KPZ}(k,t=0)\sim k^{-2\beta z-d}$ with the KPZ exponents instead, giving $n_{\hat{\bm{k}}}\propto k^{-2.77}$. This naive guess coincides with the KPZ asymptotic behavior observed in simulations of the noisy Kuramoto-Sivashinsky equation [104], as well as with predictions from acoustic turbulence in a 2d superfluid Bose gas [105]
Note that for $d=1$, both EW and KPZ universality classes produce the same low-momentum power-law decay $n_{\hat{k}}\propto 1/k^{2}$, as confirmed by numerical simulations [26], which is no longer the case in 2d.

### v3

@@ IV.7 Distribution of condensate phase and density fluctuations
where $\Delta t\in\left[35{\rm ps},350{\rm ps}\right]$ as shown on Fig. 3, is displayed on Fig. 9. The $\text{sign}(\lambda)$ in (18) stems for the negative value obtained for $\lambda$ with our parameters, which flips the distribution into its mirror image compared with the theoretical predictions obtained for $\lambda>0$ [22, 24]. The distribution of phase fluctuations exhibits clear non-Gaussian tails. As $\Delta t/t_{0}\sim 0.1$ in our simulations, we expect it to be described by the Pearson distribution. The latter is represented with a dashed blue line together with our numerical data on Fig. 9, showing a perfect agreement with the expected KPZ distribution.
@@
Figure 10: Histograms of the density $n$ normalized by its mean-field value $n_{0}$ corresponding to the EW, KPZ, and vortex points of Fig. 2. Each histogram contains approximately $5\times 10^{5}$ data. The two formers are fitted by a Gumbel distribution of parameter $m=10$, mean $\langle n\rangle/n_{0}$, and variance $\text{Var}\left[n/n_{0}\right]$, shown by the dashed blue line. The latter is fitted by an exponential distribution of parameter $p=3.5$.
First we observe that the density distribution at the EW and at the KPZ point perfectly match, which is not too surprising from Figs. 2 and 4. These distributions have asymmetric tails, stemming from the constraint $n\geq 0$. In analogy with photon lasers operating above threshold [111], one could expect the density fluctuations to have a Poisson distribution, but we find that our data is not well described by such a distribution, and instead can be modeled by a Gumbel distribution of parameter $m=10$. Density fluctuations being related to that of the phase through (1), it remains open whether or not KPZ fluctuations can leave a signature in the density. A definite conclusion requires further studies which are beyond the scope of this work. At the vortex point, the density distribution markedly differs and is rather fitted by an exponential distribution. This indicates again that the formation of defects occurs randomly, following a Poisson process. Similar exponential distributions have been reported in the nucleation time distribution of spiral defects, in the context of the 2d CGLE [112], and the discrepancies on Fig. 10 between the histogram and the exponential distribution at large $n/n_{0}$ can be explained by finite size effects [71].
@@ V Conclusion
In this paper, we have explored the universal properties of a 2d exciton-polariton condensate under incoherent pumping through numerical simulations of a generalized Gross-Pitaevskii equation. Controlling the effective nonlinearity of the phase dynamics through the interaction strength between reservoir excitons and polaritons, we have identified three different regimes: a first one for small nonlinearity where the dynamics of the density and of the phase are decoupled, and where the phase exhibits EW universal properties, a second one at intermediate nonlinearity where the density and phase are still decoupled but where the phase roughens with properties belonging to the KPZ universality class, a third one for large values of the nonlinear parameter, where density and phase fluctuations are coupled, since spatial vortices induce both phase singularities and deep minima in the condensate density at the positions of the vortex cores.
@@
We have characterized the properties of these three regimes by studying the condensate first-order correlation function, the momentum distribution, and the distribution of the phase and density fluctuations. In the weakly nonlinear limit, we report evidence of an algebraic decay of the first-order correlation function indicating the emergence of the EW universality class, from which we establish, upon increasing the interaction strength a slow crossover towards stretched exponential decay with exponents of the 2d KPZ class, and eventually a very rapid decay leading to saturation in the vortex phase. This observed behaviors in the EW and KPZ regimes are then confirmed by a careful study of finite size effects in the condensate decoherence time. The different phases are also distinguished by their momentum density distribution, where notably the EW and KPZ regimes exhibit different scaling at small momenta, in contrast with 1d condensates.
@@
Finally, we have focused on the statistical properties of phase fluctuations to provide a further validation of the three identified regimes. We have found that the distribution of phase fluctuations is Gaussian in the EW weakly nonlinear regime, it displays non-Gaussian tails at stronger non-linearity in the KPZ regime, in agreement with those obtained in numerical simulations of interfaces in the 2d KPZ class, whereas the distribution is exponential in the vortex regime. Furthermore, we observe that the probability distribution of the density fluctuations does not have a simple form – Gaussian or Poissonian – in the EW and KPZ regimes.
@@
Our analysis highlights that earlier observations from experiments [79] and numerical simulations [34] correspond to the realization of the EW regime and provides a microscopic prediction for the observed power-law decays and the ratio of the corresponding exponents. This also clarifies the distinction between closed and dissipative condensates, while showing that quasi-order can be accessed in a weakly non-linear out-of-equilibrium configuration.
@@
Our work provides a unified picture of the universal properties of 2d polariton condensates in various parameter regimes. In outlook, it would be of great interest to provide a quantitative prediction for the transition to the vortex regime, whose qualitative picture has been developed in Refs. [30, 31], as well as a deeper understanding of oscillations observed in density correlations, and the extension of the proposed phase diagram to pumping rates close to the condensation threshold where large proliferation of space-time defects is expected. It could be also interesting to further study the distributions of density fluctuations in the EW and KPZ regimes and provide a theoretical explanation of the observed statistics. Overall, our findings pave the way for further experimental explorations of the universal properties of driven-dissipative quantum systems.
@@ Acknowledgements.
We warmly thank Ivan Amelio, Marzena Szymańska, Harvey Weinberger and Vincent Rossetto for stimulating discussions. FH acknowledges support from the Laboratoire d’excellence LANEF in Grenoble (ANR-10-LABX-51-01), AM acknowledges funding from the Quantum-SOPHA ANR Project ANR-21-CE47-0009. LC acknowledges support from Institut Universitaire de France (IUF). This work was partly supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project ARQADIA, grant agreement no. 949730), and under Horizon Europe research and innovation programme (ANAPOLIS, grant agreement no. 101054448).
@@ Appendix A Adiabatic Bogoliubov theory
In the adiabatic approximation of the reservoir $\partial_{t}n_{R}\approx 0$, Eq. (20) reduces to $n_{R}=P/(\gamma_{R}+R|\psi|^{2})$. The condensate wavefunction is expanded in fluctuations $\delta\psi$ around its stationary mean field density $n_{0}=\gamma_{R}(p-1)/R$, as $\psi(\bm{r},t)=e^{-i{\color[rgb]{0,0,0}\mu_{\rm th}}t}\left[\sqrt{n_{0}}+\delta\psi(\bm{r},t)\right]$. Injecting it in Eq. (19) leads at zeroth order to the mean field exciton reservoir density and condensate blueshift, respectively $n_{R0}=\gamma_{0}/R$ and ${\color[rgb]{0,0,0}\mu_{\rm th}}=gn_{0}+2g_{R}n_{R0}$. To first order in $\delta\psi$, Eq. (19) reduces to the linear dynamics
@@
where we introduced $\tilde{\xi}=\xi e^{-i{\color[rgb]{0,0,0}\mu_{\rm th}}t}$ and the Bogoliubov matrix
@@
[…]
@@
$\omega_{\pm}=-i\Gamma_{\hat{\bm{k}}}\pm\sqrt{E_{\hat{\bm{k}}}^{2}-g_{i}^{2}n_{0}^{2}}\,$, (27)
@@ A.1 Phase-phase correlator
$+\dfrac{\langle\delta n_{\bm{k},\omega}\delta n_{-\bm{k},-\omega}\rangle}{4n_{0}^{2}},$
${\color[rgb]{0,0,0}\langle\delta n_{\bm{k},\omega}\delta n_{-\bm{k},-\omega}\rangle}$ $=$ ${\color[rgb]{0,0,0}\dfrac{\sigma n_{0}(\epsilon_{\bm{\hat{k}}}^{2}+\omega^{2}+(\Gamma_{\bm{\hat{k}}}+g_{i}n_{0})^{2})}{(\omega^{2}-\omega_{-}^{2})(\omega^{2}-\omega_{+}^{2})}.}$
The large scale properties of the condensate can be understood taking advantage of the splitting between these two branches at the crossover momentum $k_{c}=\sqrt{2n_{0}|m|}\left[\sqrt{g_{e}^{2}+g_{i}^{2}}-|g_{e}|\right]^{1/2}$, separating the infrared collective modes $k\ll k_{c}$ from the ultraviolet free particles $k\gg k_{c}$. Focusing on the former limit and choosing $\epsilon_{\hat{\bm{k}}}=k^{2}/2m$ and $\gamma_{\ell}(\hat{\bm{k}})=\gamma_{0}+\gamma_{2}k^{2}$, we find for $k\approx 0$ that $\omega_{-}\approx-2ig_{i}n_{0}$ and $\omega_{+}\approx-i\nu k^{2}$, where $\nu$ is the effective phase viscosity (4) introduced in the main text.
@@
$\times\dfrac{k^{2}}{\left(\omega^{2}+4g_{i}^{2}n_{0}^{2}\right)\left(\omega^{2}+\nu^{2}k^{4}\right)}$
where $C_{nn}(r,t)=\langle\delta n(r,t)\delta n(0,0)\rangle-\langle\delta n(0,0)^{2}\rangle$. The first term, which is the Edwards-Wilkinson two-point correlation defined by (30), dominates the scaling behavior of $C_{\theta\theta}(r,t)$ in the infrared regime. This can be established by integrating over the frequency, and analyzing the small $k$ behavior of the integrals at fixed $t$ and $r$. One finds that, for all the terms but $C_{\rm EW}(r,t)$, the integrals are convergent in the IR, hence they vanish at large $r$. They are dominated by the UV cut-off $\Lambda$. In contrast, in the EW contribution, the leading term at small $k$ diverges in the IR as $\sim\int dk/k$ for $d=2$ and thus dominates the IR behavior, as is detailed below.
The EW correlation reads [2, 45]
@@
where $\mathcal{C}_{0}=2\frac{D}{\nu}(4\nu)^{2\beta}$ and $y_{0}=2\sqrt{\nu}$ are non-universal constants, with $D$ and $\nu$ the effective parameters defined in Eqs. (4) and (6). This function is characterized by the universal dynamical exponent $z=2$, the growth exponent $\beta=\frac{2-d}{4}$, and the universal scaling function [113, 45].
$F^{(d)}(y)=\int\frac{d\bm{\tilde{k}}}{(2\pi)^{d}}\frac{1-e^{i\bm{\tilde{k}}\cdot\bm{y}-\bm{\tilde{k}}^{2}/4}}{\bm{\tilde{k}}^{2}}\quad\hbox{with}\quad\bm{\tilde{k}}=2\sqrt{\nu t}\bm{k}$ (31)
Let us analyze the finite-size effects on $C_{\rm EW}(r=0,t)$. For a system of linear size $L$, from the diffusive nature of EW modes, one expects a change in its behavior at a time $t^{*}\sim L^{2}/\nu$. Adapting the arguments given in Ref. [92], at fixed $t$, one can split the dimensionless momentum integral of $F^{(d)}(y)$ into two: one for ${k}\leq 1/\sqrt{\nu t}$, the other for $k>1/\sqrt{\nu t}$.
In the limit of small times $t\ll t^{*}$, the bound $1/\sqrt{\nu t}$ is large, such that both integrals contain enough modes to be treated as continuous integrals. In this case, in dimension $d=1$, the integrals can be performed analytically, yielding $C_{\rm EW}(0,t\ll t^{*})=\dfrac{2D}{\sqrt{\pi\nu}}\sqrt{t}$ consistent with the scale invariant form (29). In dimension $d=2$, one can expand to second order $e^{-\tilde{k}^{2}/4}$ in the first integral and send the UV cutoff $\Lambda$ to infinity in the convergent part $\propto e^{-\tilde{k}^{2}/4}$ of the second integral. Both contributions are finite, and the result is thus dominated in the IR by the $\sim\int dk/k$ contribution, leading to $C_{\rm EW}(0,t)\approx\dfrac{D}{2\pi\nu}\ln(\nu t\Lambda^{2})$, i.e. a logarithmic time dependence [65, 66, 114, 45, 56] with $\Lambda$ a momentum cutoff.
In the limit of large times $t\gg t^{*}$, the bound $1/\sqrt{\nu t}$ tends to zero. In this case, the first integral should be replaced by the discrete sum over the lowest modes. Expanding $e^{-\tilde{k}^{2}/4}$, one obtains that in both $d=1$ and 2, this term yields the dominant contribution at large time which is proportional to $t$. In $d=1$, it reads $C_{\rm EW}(0,t)\approx\dfrac{2D}{\mathcal{L}}t$ where $\mathcal{L}$ is a length scale maintaining homogeneity. A natural choice for this length scale, stemming from the lowest momenta of the finite-size system, is $\mathcal{L}=\uptau^{(1)}L$, with $\uptau^{(1)}$ a dimensionless constant [97]. This constant contains the dependence of the condensate linewidth on the microscopic parameters that goes beyond the Bogoliubov approximation. In $d=2$, one obtains $C_{\rm EW}(0,t)\approx\dfrac{D}{\uptau^{(2)}L^{2}}t$ with the dimensionless constant $\uptau^{(2)}$.
@@
From this finite-size study, we expect $g^{(1)}(t\gg t^{*})\sim e^{-t/\uptau_{c}(L)}$ with the coherence time $\uptau_{c}(L)\propto L^{z-2\beta z}$, generalizing the result of [115, 97] to 2d systems. This Schawlow-Townes-like limit [98], following EW or KPZ features, means that the growing phase interface looses the memory of its initial condition at long times and behaves as a Brownian motion in time [99].
@@ A.2 Momentum distribution
$n_{\hat{\bm{k}}}=(2\pi)^{d}n_{0}\delta(\bm{k})+\dfrac{\sigma}{\Gamma_{\hat{\bm{k}}}}\left[1+n_{0}^{2}\dfrac{g_{e}^{2}+g_{i}^{2}}{E_{\hat{\bm{k}}}^{2}+\Gamma_{\hat{\bm{k}}}^{2}-g_{i}^{2}n_{0}^{2}}\right].$ (33)
Replacing $\epsilon_{\hat{\bm{k}}}=k^{2}/2m$ and $\gamma_{\ell}(\hat{\bm{k}})=\gamma_{0}+\gamma_{2}k^{2}$, we obtain $n_{\hat{\bm{k}}}\underset{k\to+\infty}{\sim}\sigma/\gamma_{2}k^{2}$ in the ultraviolet regime. The short-momentum behavior is given, up to a constant, by $n_{\hat{\bm{k}}}\underset{k\to 0^{+}}{\sim}\tilde{C}_{\rm EW}(k,t=0)$, where $\tilde{C}_{\rm EW}(k,t)=\int d\bm{r}e^{-i\bm{k}\cdot\bm{r}}C_{\rm EW}(r,t)$.
At equal time, we obtain $\tilde{C}_{\rm EW}(k,0)\sim k^{-2\beta z-d}$. From this, we deduce that in the EW regime $n_{\hat{\bm{k}}}\propto k^{-2}$. This tail in the momentum distribution is similar to that of a thermal closed Bose gas, from which an effective polariton temperature can be defined as discussed in the main text. Since this result was obtained using the linear theory, we can expect the non-linearity of the phase dynamics to affect this regime and rather lead to $n_{\hat{\bm{k}}}\underset{k\to 0^{+}}{\sim}\tilde{C}_{\rm KPZ}(k,t=0)\sim k^{-2\beta z-d}$ with the KPZ exponents instead, giving $n_{\hat{\bm{k}}}\propto k^{-2.77}$. This naive guess coincides with the KPZ asymptotic behavior observed in simulations of the noisy Kuramoto-Sivashinsky equation [116], as well as with predictions from acoustic turbulence in a 2d superfluid Bose gas [117]. This departure from the thermal-like tail $n_{\hat{\bm{k}}}\propto k^{-2}$ can be interpreted as a signature of the purely non-equilibrium nature of the KPZ equation in 2d.
Note that for $d=1$, both EW and KPZ universality classes produce the same low-momentum power-law decay $n_{\hat{k}}\propto 1/k^{2}$, as confirmed by numerical simulations [26]. This is a consequence of the accidental time-reversal symmetry of the KPZ equation in $d=1$, which yields a fluctuation-dissipation relation and leads to the same equal-time behavior as the EW equation, although the temporal beahvior is distinct. This time-reversal symmetry no longer holds in $d\neq 1$ for the KPZ equation [59].